Determine the magnitude P of the force which the man must exert perpendicular to the handle of the high – pressure washer in order to cause loss of contact at the front support B. note that the operator prevents movement of the wheel with his left foot. The 60-Kg machine has its mass center at point G. Treat the problem as two – dimensional.  

 



1
Expert's answer
2021-07-26T17:11:02-0400

"m=60kg"

"W=mg"

For water to loss contact at point B, net torque about point 'O' must be zero i.e


"\sum M_o=0"

"=P(800)-W(125)+R_b(310+125)=0"


When washer is about to loss contact at B, "R_B=0"


So;"P=\frac {60×9.61×125}{800}"


"P=91.96875N"